# AnyEnhance: A Unified Generative Model with Prompt-Guidance and Self-Critic for Voice Enhancement

### v1

@@ V-E2 Effect of Prompt Guidance
The results in Table IV were obtained without using any prompt information during inference (except for TSE Group). In this section, we explore the effect of introducing prompt guidance by leveraging audio samples from the same speaker as prompts during inference on the Librivox GSR, CCMusic GSR, and Voicefixer SR datasets. As shown in Table VII, incorporating prompt information leads to further improvements in performance across all metrics, and among all datasets. These results indicate that incorporating speaker-specific prompt information during inference can effectively enhance the restoration quality of the model, further demonstrating the adaptability of AnyEnhance under speaker-aware scenarios.
V-E3 Effect of Self-Critic Sampling
(a) OVRL Comparison
(b) SpeechBERTScore Comparison
(c) Similarity Comparison
Figure 3: Comparison between Self-Critic and non-Self-Critic. Using Self-Critic Sampling consistently achieves better performance across varying sampling timesteps and metrics.
@@ V-E3 Effect of Self-Critic Sampling
In this ablation study, we investigate the effect of the critic head on the model performance. We evaluate the model with and without the critic head on the Librivox GSR dataset across different inference steps. As shown in Figure 3, the model with the critic head consistently outperforms the version without it across different inference steps, in terms of OVRL, SpeechBERTScore, and Similarity metrics. This demonstrates the importance of the critic head in guiding the model to generate high-quality audio outputs.
@@
V-E4 Effect of Improved Data Simulation Process
In this ablation study, we evaluate the effect of the improved data simulation process on the model performance. We train the same model on the same dataset with and without the improved data simulation process described in Section III-A, and evaluate the model on the Librivox GSR and Voicefixer GSR datasets. The Librivox GSR dataset is simulated with the improved data simulation process, while the Voicefixer GSR dataset is simulated using the old-fasioned data simulation process as introduced in Section III-A. The results are shown in Table VIII. Even though the Voicefixer GSR dataset is not simulated with the improved data simulation process, the model trained with the improved data simulation process still outperforms the model trained without it across all metrics (except BAK on the Voicefixer GSR dataset), demonstrating the effectiveness of the improved data simulation process in enhancing generalization and robustness of AnyEnhance.
TABLE VIII: Effect of Improved Data Simulation Process, w/o improved data simulation means the model is trained using $d_{\text{noise}}(d_{\text{bw}}(d_{\text{clip}}(d_{\text{reverb}}(\cdot))))$ described in Section III-A, w/ improved data simulation means the model is trained using $d_{\text{bw}}(d_{\text{clip}}(d_{\text{reverb}}(d_{\text{noise}}(d_{\text{vocal\_effect}}(\cdot)))))$ in addition to traditional data simulation.
@@ V-E5 Effect of One Model for All Tasks
TABLE IX: Effect of one model for all tasks, we compare models trained on each single task’s distortion against a unified model trained on all tasks simultaneously. The unified multitask training approach consistently outperforms single-task model across all datasets.
@@
In this ablation study, we compare the performance of AnyEnhance trained individually on each task against a unified model trained on all tasks simultaneously. The results are shown in Table IX. The results show that the unified model outperforms task-specific models across most metrics. For instance, on the Librivox GSR dataset, the unified model achieves higher SIG (3.504 vs. 3.438) and NISQA (4.191 vs. 3.827) scores compared to the task-specific model. Similar trends are observed in other tasks, such as SE, SR, and TSE, where the unified model achieves consistently higher OVRL and NISQA scores.
@@ VI Conclusion
In this work, we introduce AnyEnhance, a unified voice enhancement model for both speech and singing voice. Using a masked generative model as the base, AnyEnhance is capable of addressing diverse enhancement tasks, including speech denoising, dereverberation, super-resolution, and target speaker extraction. We incorporate a prompt-guidance mechanism for in-context learning, which allows the model to natively accept a reference speaker’s timbre without altering the underlying architecture. This strategy boosts enhancement performance when a reference audio is available and enable the target speaker extraction task without altering the underlying architecture. Moreover, we integrate a self-critic mechanism into the generative process for masked generative models, yielding higher-quality outputs through iterative self-assessment and refinement. Extensive experiments on various enhancement tasks demonstrate that AnyEnhance outperforms existing methods in terms of both objective metrics and subjective listening tests.

### v3

@@ V-E2 Effect of Prompt Guidance and Self-Critic Sampling
In this section, we investigate the effect of prompt guidance and self-critic sampling on the performance of AnyEnhance. The results in Table IV were obtained without using any prompt information during inference (except for TSE Group). We first evaluate the performance of AnyEnhance with and without prompt guidance on the Librivox GSR, CCMusic GSR, and Voicefixer SR datasets. The results are shown in Table X. The model with prompt guidance consistently outperforms the model without prompt guidance across all datasets and metrics (except for BAK on Voicefixer SR). These results indicate that incorporating speaker-specific prompt information during inference can effectively enhance the restoration quality of the model, further demonstrating the adaptability of AnyEnhance under speaker-aware scenarios.
(a) Librivox GSR - OVRL
(b) Librivox GSR - SpeechBERTScore
(c) Librivox GSR - Similarity
(d) CCMusic SR - OVRL
(e) CCMusic SR - SpeechBERTScore
(f) CCMusic SR - Similarity
(g) VCTK TSE - OVRL
(h) VCTK TSE - SpeechBERTScore
(i) VCTK TSE - Similarity
Figure 4: Comparison between Self-Critic and non-Self-Critic across three tasks: Librivox GSR, CCMusic SR, and VCTK TSE. Using Self-Critic Sampling consistently leads to better performance across varying sampling timesteps and metrics (OVRL, SpeechBERTScore, and Speaker Similarity).
(a) OVRL
(b) SpeechBERTScore
(c) Speaker Similarity
Figure 5: Comparison of the combined and individual effects of Prompt-Guidance and Self-Critic Sampling across different inference steps. Both mechanisms contribute to performance improvements, with prompt-guidance generally offering a larger gain.
@@
To investigate the effectiveness of the self-critic sampling mechanism, we evaluate the model with and without the critic head on 3 different datasets: Librivox GSR, CCMusic SR, and VCTK TSE across different inference steps. The results are shown in Figure 4. We observe that the model with the critic head consistently outperforms the model without it across different inference steps, in terms of OVRL, SpeechBERTScore, and Similarity metrics. This demonstrates the importance of the critic head in guiding the model to generate high-quality audio outputs.
@@
Finally, we evaluate the combined effect of prompt guidance and self-critic sampling on the model performance. The results are shown in Figure 5. We observe that both mechanisms contribute to performance improvements, and prompt guidance generally offering a larger gain. This demonstrates that the two mechanisms can be synergistically combined to enhance the model performance.
V-E3 Effect of Singing-Voice-oriented Design
In this section, we conduct an ablation study to evaluate the effect of the singing-voice-oriented design on the model performance, namely the improved data simulation process and the inclusion of singing voice data. First we train the same model on the same dataset with and without the improved data simulation process described in Section III-A, and evaluate the model on the Librivox GSR and Voicefixer GSR datasets. The Librivox GSR dataset is simulated with the improved data simulation process, while the Voicefixer GSR dataset is simulated using the old-fasioned data simulation process as introduced in Section III-A. The results are shown in Table XI. Even though the Voicefixer GSR dataset is not simulated with the improved data simulation process, the model trained with the improved data simulation process still outperforms the model trained without it across all metrics (except BAK on the Voicefixer GSR dataset), demonstrating the effectiveness of the improved data simulation process in enhancing generalization and robustness of AnyEnhance.
To better understand the singing-voice-oriented design, we conduct an ablation study to evaluate both the improved data simulation process and the inclusion of singing voice data on the model performance. We train the model with and without the improved data simulation process and the inclusion of fullband singing voice dataset. The results are shown in Table XII. Both the improved data simulation process and the inclusion of singing voice data contribute to the performance improvements across all metrics, with the improved data simulation process introduces gain on the overall quality (OVRL, NISQA), and the inclusion of singing voice data improves largely on the performance on SpeechBERTScore and speaker similarity. The model trained with both enhancements achieves the best performance, demonstrating the effectiveness of these design choices in enhancing the model’s ability to handle singing voice scenarios.
TABLE XI: Effect of Improved Data Simulation Process, w/o improved data simulation means the model is trained using $d_{\text{noise}}(d_{\text{bw}}(d_{\text{clip}}(d_{\text{reverb}}(\cdot))))$ described in Section III-A, w/ improved data simulation means the model is trained using $d_{\text{bw}}(d_{\text{clip}}(d_{\text{reverb}}(d_{\text{noise}}(d_{\text{vocal\_effect}}(\cdot)))))$ in addition to traditional data simulation.
@@ V-E3 Effect of Singing-Voice-oriented Design
TABLE XII: Ablation study on singing-voice-oriented design. We examine the effects of (1) improving the data simulation process and (2) including singing voice data. Results show improvements across all evaluation metrics, particularly in SpeechBERTScore and speaker similarity.
[…]
CCMusic GSR 2.981 3.736 2.542 2.916 0.559 0.776
3.17 3.77 2.75 3.406 0.598 0.839
3.144 3.537 2.694 3.124 0.739 0.863
3.213 3.647 2.809 3.358 0.759 0.889
V-E4 Effect of REPA module and Semantic Feature
TABLE XIII: Ablation study on REPA module and semantic features across Librivox GSR and Voicefixer GSR test sets. Both REPA and semantic features contribute to performance improvements.
[…]
Librivox GSR AnyEnhance 3.505 4.135 3.256 4.200 0.763 0.913
- REPA 3.481 4.141 3.242 4.082 0.764 0.923
      - Noisy Emb 3.430 4.123 3.178 3.975 0.734 0.902
Voicefixer GSR AnyEnhance 3.400 4.022 3.103 4.329 0.831 0.907
- REPA 3.372 4.024 3.082 4.210 0.830 0.915
      - Noisy Emb 3.343 4.051 3.061 4.215 0.812 0.905
We examine the effect of the REPA module and semantic features on both the Librivox GSR and Voicefixer GSR test sets, as shown in Table XIII. Removing the REPA module leads to slight degradations in SIG, OVRL, and NISQA, indicating that REPA contributes to perceptual and quality improvements. Further removing the noisy semantic features results in consistent drops across all metrics, especially SpeechBERTScore and Similarity, highlighting the effectiveness of the semantic features in enhancing the model’s ability.
@@ V-E5 Effect of One Model for All Tasks
TABLE XIV: Effect of one model for all tasks, we compare models trained on each single task’s distortion against a unified model trained on all tasks simultaneously. The unified multitask training approach consistently outperforms single-task model across all datasets.
@@
In this ablation study, we compare the performance of AnyEnhance trained individually on each task against a unified model trained on all tasks simultaneously. The results are shown in Table XIV. The results show that the unified model outperforms task-specific models across most metrics. For instance, on the Librivox GSR dataset, the unified model achieves higher SIG (3.504 vs. 3.438) and NISQA (4.191 vs. 3.827) scores compared to the task-specific model. Similar trends are observed in other tasks, such as SE, SR, and TSE, where the unified model achieves consistently higher OVRL and NISQA scores.
@@ VI Conclusion
In this work, we introduce AnyEnhance, a unified voice enhancement model for both speech and singing voice. Using a masked generative model as the base, AnyEnhance is capable of addressing diverse enhancement tasks, including speech denoising, dereverberation, super-resolution, and target speaker extraction. We incorporate a prompt-guidance mechanism for in-context learning, which allows the model to natively accept a reference speaker’s timbre without altering the underlying architecture. This strategy boosts enhancement performance when a reference audio is available and enable the target speaker extraction task without altering the underlying architecture. Moreover, we integrate a self-critic mechanism into the generative process for masked generative models, yielding higher-quality outputs through iterative self-assessment and refinement. Extensive experiments on various enhancement tasks demonstrate that AnyEnhance outperforms existing methods in terms of both objective metrics and subjective listening tests. Future work includes constructing more realistic degradation simulation scenarios, exploring textual prompts for guided enhancement, and learning optimal task-specific loss weights for multitask training.
Acknowledgments
This work is funded by the NSFC under Grant 62376237, Shenzhen Science and Technology Program ZDSYS20230626091302006, 2023 Shenzhen stability Science Program, and Internal Project Fund from Shenzhen Research Institute of Big Data (Grant No. T00120230002).